OpenStudy (anonymous):

how do you convert 2.5lb to oz

OpenStudy (anonymous):

2.5*16

OpenStudy (anonymous):

How many ounces in one pound?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

\[\large =2.5lbs \times \frac{16 oz}{lb}\]

OpenStudy (anonymous):

well it be 0.2oz

OpenStudy (anonymous):

nope. it's 40 oz.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

to convert pounds to ounces, u multiply the amount u need to convert by 16 to get the number of ounces. so it's 2.5*16

OpenStudy (anonymous):

do use the 1 pound too

OpenStudy (anonymous):

\[\large 2.5 \cancel{pounds} \times \frac{16 \space ounces}{\cancel{1 pound}}\]

OpenStudy (anonymous):

exactly what i just saud

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes. It shows how the units cancel, though. @jalyn1 asked what happens to the pounds.

OpenStudy (anonymous):

thankz